# VITA: Vision-to-Action Flow Matching Policy

### v2

@@ Appendix F Task Settings
F.1 AV-ALOHA Simulation tasks.
@@
F.2 AV-ALOHA Real tasks
@@
F.3 Robomimic Tasks
Robomimic is a benchmark of single-arm imitation learning tasks [24]. We adapt the environment for compatibility with the LeRobot [38] codebase. The robot state includes arm joint positions (encoded with $\sin$ and $\cos$), joint velocities, end-effector pose, gripper finger positions, and gripper finger velocities. The action space consists of six values for delta position control of the end-effector pose and one value for the absolute position of the gripper.
@@ F.3 Robomimic Tasks
Figure 12: Hyperparameter ablations on ThreadNeedle. Each sub-plot varies a single coefficient while holding all others at their default values in  Table˜6. Top: action encoder reconstruction weight, flow latent decoding weight, action encoder contrastive weight. Bottom: latent flow contrastive weight, latent flow consistency weight, action KL weight.
Appendix G Training
@@
In each plot of  Figure˜12, we sweep a single weight while holding all other hyperparameters at the default ( Table˜6), and visualize the success rates over training steps. With these experimental results, a robust configuration on ThreadNeedle includes moderate weights for both AE and FLD (typically in $[0.5,1.0]$), moderate FLC, minimal or no contrastive penalties, and no KL regularization on the action and vision latent.
Appendix H Implementations
H.1 VITA
@@
VITA encodes observations into a latent vector $\bm{z}_{0}$ using a ResNet-18 backbone. The flow network, which learns the mapping from $\bm{z}_{0}$ to $\hat{\bm{z}}_{1}$, is implemented as an MLP. We use a 6-step Euler ODE solver. The latent action $\hat{\bm{z}}_{1}$ is translated to action chunks by a lightweight MLP-based action decoder. A summary of all hyperparameters and loss weights is provided in Table˜6.
H.2 Flow Matching Policy
@@
The FM policy learns a velocity field using a transformer backbone, with AdaLN for conditioning. A 6-step Euler solver is used for solving ODEs. The hyperparameters and loss weights are summarized in Table˜7.
H.3 Diffusion Policy
@@
We follow the DP  [10] implementation in LeRobot [38]. The core of the policy is a U-Net noise predictor, which uses FiLM [31] for conditioning. The model is trained for 100 timesteps using a cosine beta schedule. During inference, trajectories are generated using a 10-step DDPM sampler. All hyperparameters and loss weights are listed in Table˜8.
H.4 Action Chunking Transformer
@@
We follow the ACT  [26] implementation in LeRobot [38]. ACT is a conditional variational autoencoder (cVAE) that generates action chunks conditioned on vision. Its architecture consists of a vision encoder for processing observations and a transformer-based decoder that models the distribution of future actions conditioned on the visual input. A complete list of hyperparameters and loss weights is provided in Table˜9.
@@
Inference steps null (defaults to training timesteps)

### v3

@@ Appendix D Task Settings
PushT 2 2 20 $\times96$ image
CloseBox 9 9 20 $\times256$ head_cam
@@
D.1 AV-ALOHA Simulation tasks.
@@
D.2 ALOHA Real Tasks
Figure 18: Autonomous rollouts of VITA on two challenging real-world AV-ALOHA tasks, HiddenPick, and TransferFromBox.
D.2.1 Bimanual Manipulation with Active Vision
To evaluate the effectiveness of VITA in real-world settings, we deploy the policy on two challenging bimanual manipulation tasks using AV-ALOHA [chuang2024active]. Both tasks are long-horizon and contain multiple stages to succeed. Examples of autonomous rollouts are shown in Figure˜18. Both tasks require precise and coordinated control of three arms, including one arm carrying an active vision camera, and two arms for bimanual manipulation.
@@ D.2.1 Bimanual Manipulation with Active Vision
D.2.2 Single-Arm Manipulation Tasks
Each single-arm ALOHA task consists of multiple stages, and includes substantial environment randomization as detailed below.
PickBall: Pick up a ball, then place it into the box (50 episodes). Both the ball and the target box appear in varying positions.
ToothBrush: Pick up the toothbrush from the side slot of the toothbrush cup, lift it up, and place it into the toothbrush cup (50 episodes). The cup location and the orientation of the side slot are randomized.
StoreDrawer: Pick up the object, put it in the drawer, and close the drawer (100 episodes). We randomize the shapes and colors of the objects as well as the positions and partial openings of the drawers. We also evaluate the out-of-distribution success rates for unseen combinations of object colors and shapes (see Appendix  E.2).
D.3 Robomimic Tasks
Robomimic is a benchmark of single-arm imitation learning tasks [robomimic2021]. We adapt the environment for compatibility with the LeRobot [cadene2024lerobot] codebase. The robot state includes arm joint positions (encoded with $\sin$ and $\cos$), joint velocities, end-effector pose, gripper finger positions, and gripper finger velocities. The action space consists of six values for delta position control of the end-effector pose and one value for the absolute position of the gripper.
@@ D.3 Robomimic Tasks
D.4 Other Tasks
PushT: Push a 2D T-shaped object into a matching T-shaped target region on the plane; achieving a coverage score >0.95>0.95 is deemed a success (200 episodes).
CloseBox: Close the lid of a paper box by manipulating the flap (200 episodes).
Appendix E Additional Experiment Results
E.1 Robustness to Online Perturbations
As depicted in Figure˜19, VITA demonstrates strong robustness to online perturbations during real-time control.
Figure 19: Robustness to online perturbations during inference on PickBall. We manually move the ball multiple times before the pick, and move the box multiple times after the pick. In both cases, the arm continues to adjust in real time and successfully reaches the correct ball and box positions.
E.2 Generalization to Unseen Objects
As shown in Figure˜20, we train the StoreDrawer task using 7 objects. Table˜3 shows that VITA achieves the highest success rate on these in-distribution objects. To further evaluate generalization, we introduce four unseen test objects with novel geometries, such as a triangular prism and a star-shaped block. On this out-of-distribution (OOD) set, VITA succeeds in picking all objects and storing them in the drawer.
Table 8: OOD SRs on four unseen objects for the StoreDrawer task.
VITA DP FM ACT
4/4 4/4 3/4 2/4
Figure 20: OOD evaluation on StoreDrawer. Training uses 8 in-distribution object–color combinations (left). Evaluation uses four unseen objects with novel shapes (right), including a triangular prism and a star-shaped block.
Appendix F Training
@@
Figure 21: Hyperparameter ablations on ThreadNeedle. Each sub-plot varies a single coefficient while holding all others at their default values in Table˜9. Top: action encoder reconstruction weight, flow latent decoding weight, action encoder contrastive weight. Bottom: latent flow contrastive weight, latent flow consistency weight, action KL weight.
@@
In each plot of Figure˜21, we tune a single hyperparameter while holding all other hyperparameters at the default (see Table˜9), and visualize the success rates over training steps. With these experimental results, a robust configuration on ThreadNeedle includes moderate weights for both AE and FLD (typically in $[0.5,1.0]$), moderate FLC, minimal contrastive penalties, and no KL regularization on action and visual latents.
Appendix G Implementations
G.1 VITA
@@
VITA encodes observations into a latent vector $\bm{z}_{0}$ using a ResNet-18 backbone. The flow network, which learns the mapping from $\bm{z}_{0}$ to $\hat{\bm{z}}_{1}$, is implemented as an MLP. We use a 6-step Euler ODE solver. The latent action $\hat{\bm{z}}_{1}$ is translated to action chunks by a lightweight MLP-based action decoder. A summary of all hyperparameters and loss weights is provided in Table˜9.
G.2 Flow Matching Policy
@@
The FM policy learns a velocity field using a transformer backbone, with AdaLN for conditioning. A 6-step Euler solver is used for solving ODEs. The hyperparameters and loss weights are summarized in Table˜10.
G.3 Diffusion Policy
@@
We follow the DP  [chi2023diffusion] implementation in LeRobot [cadene2024lerobot]. The core of the policy is a U-Net noise predictor, which uses FiLM [perez2018film] for conditioning. The model is trained for 100 timesteps using a cosine beta schedule. During inference, trajectories are generated using a 10-step DDPM sampler. All hyperparameters and loss weights are listed in Table˜11.
G.4 Action Chunking Transformer
@@
We follow the ACT  [zhao2023learning] implementation in LeRobot [cadene2024lerobot]. ACT is a conditional variational autoencoder (cVAE) that generates action chunks conditioned on vision. Its architecture consists of a vision encoder for processing observations and a transformer-based decoder that models the distribution of future actions conditioned on the visual input. A complete list of hyperparameters and loss weights is provided in Table˜12.
@@
Inference steps DDPM 10